Evaluate the integral of (5x+2)³ dx


1
Expert's answer
2021-03-24T15:03:32-0400


Answer:

"\int (5x+2)^3dx=\dfrac{1}{20}(5x+ 2)^4+C"


Explanation:

"\int (5x+2)^3dx ="

Making the substitution,

"v =5x+2", "\dfrac{dv}{dx} =5"


"dx = \dfrac{dv}5"


then,

"\int{(5x+2)^3}dx = \int v³\ \dfrac{dv}{5} = \dfrac15\int v³\ dv"


"= \dfrac15(\dfrac{v⁴}{4}) + C = \dfrac{v⁴}{20} + C"


Since, "v = 5x + 2"


"\therefore \int (5x+2)^3dx=\dfrac{1}{20}(5x+ 2)^4+C"